# Decoding Futures Price Dynamics: A Regularized Sparse Autoencoder for Interpretable Multi-Horizon Forecasting and Factor Discovery

### v3

@@ 5.1 Experimental Setup
  • Up-Move Hit Rate (UHR) & Down-Move Hit Rate (DHR): Calculated on the subset of days with actual price changes ($|p_{t+\tau_{i}}-p_{t}|>\epsilon_{\text{zero}}$). UHR is the proportion of actual up-moves correctly predicted as up-moves. DHR is the proportion of actual down-moves correctly predicted as down-moves.
  • Matthews Correlation Coefficient (MCC): A robust measure for binary classification quality, considering true/false positives and negatives, calculated on the up, down predictions for days with actual price changes ($|p_{t+\tau_{i}}-p_{t}|>\epsilon_{\text{zero}}$). Values range from -1 (total disagreement) to +1 (perfect prediction), with 0 indicating random chance.
@@
5.2 Predictive Accuracy Results
Performance on the test set (2021–2023) is in Table 3 and Table 4.
Table 3: Predictive Performance for Copper Log Prices (Test Set: 2021–2023).
Model Horizon RMSE MAE DA (%)
RSAE 1-day 0.0125 0.0093 58.2
5-day 0.0283 0.0210 55.1
22-day 0.0591 0.0452 53.5
Persistence 1-day $0.0152$ $0.0115$ 50.0
5-day $0.0340$ $0.0258$ 50.0
22-day $0.0705$ $0.0540$ 50.0
ARIMA 1-day $0.0149$ $0.0112$ 51.5
5-day $0.0335$ $0.0251$ 50.8
22-day $0.0690$ $0.0525$ 50.3
XGBoost 1-day $0.0135$ $0.0101$ 56.5
5-day $0.0305$ $0.0228$ 54.0
22-day $0.0630$ $0.0481$ 52.1
LSTM 1-day $0.0130$ $0.0098$ 57.3
5-day $0.0295$ $0.0220$ 54.5
22-day $0.0615$ $0.0469$ 52.8
Table 4: Predictive Performance for WTI Crude Oil Log Prices (Test Set: 2021–2023).
Model Horizon RMSE MAE DA (%)
RSAE 1-day 0.0185 0.0139 59.1
5-day 0.0410 0.0315 56.3
22-day 0.0835 0.0650 54.2
Persistence 1-day $0.0225$ $0.0170$ 50.0
5-day $0.0500$ $0.0385$ 50.0
22-day $0.1020$ $0.0805$ 50.0
ARIMA 1-day $0.0220$ $0.0165$ 51.2
5-day $0.0490$ $0.0375$ 50.6
22-day $0.1000$ $0.0780$ 50.2
XGBoost 1-day $0.0198$ $0.0150$ 57.5
5-day $0.0445$ $0.0340$ 55.1
22-day $0.0890$ $0.0695$ 52.9
LSTM 1-day $0.0192$ $0.0145$ 58.3
5-day $0.0430$ $0.0328$ 55.7
22-day $0.0865$ $0.0675$ 53.5
To provide a more nuanced view of directional forecasting skill, Table 5 presents advanced directional accuracy metrics for the 1-day horizon for Copper futures, comparing RSAE with the LSTM baseline. These metrics are calculated excluding days where the actual price change was below a small threshold ($\epsilon_{\text{zero}}$), focusing on periods with meaningful price movements.
Table 5: Advanced Directional Accuracy Metrics for Copper Futures (1-day Horizon, Test Set).
Metric RSAE LSTM
Standard DA (%) 58.20 57.30
DA Excl. No-Change (%) 59.53 58.12
Up-Move Hit Rate (%) 60.15 58.90
Down-Move Hit Rate (%) 58.88 57.31
Matthews Correlation Coefficient (MCC) 0.191 0.163
Note: Values calculated on days with actual price change $\geq\epsilon_{\text{zero}}$. Standard DA includes no-change days.
Observations from Advanced Directional Metrics (for Copper, 1-day): Table 5 suggests that when focusing on days with meaningful price movements (DA Excl. No-Change), the RSAE (59.53%) maintains a slight edge over the LSTM (58.12%). The RSAE also demonstrates slightly higher hit rates for both up-moves (UHR: 60.15% vs. 58.90%) and down-moves (DHR: 58.88% vs. 57.31%). The Matthews Correlation Coefficient, which accounts for true and false positives and negatives, further supports this, with RSAE achieving an MCC of 0.191 compared to LSTM’s 0.163. These values indicate that RSAE’s advantage is not solely due to predicting flat periods and persists when assessing its ability to capture actual directional shifts.
Observations from Accuracy Results: RSAE generally achieves competitive or superior predictive performance. RMSE/MAE values are consistently lowest. DA is strong. Performance degrades with horizon. Average $L_{0}$ norm for $\bm{z}^{*}$ was $\approx 5.8$ (Copper) and $\approx 6.3$ (Crude Oil).
[…]
Table 6: Diebold-Mariano Test p-values (One-sided: $H_{1}$: RSAE better than LSTM).
Horizon Copper Futures WTI Crude Oil Futures
1-day 0.028 0.035
5-day 0.045 0.061
22-day 0.053 0.623
The DM test results (Table 6) indicate that for 1-day ahead forecasts, RSAE significantly outperforms the LSTM model for both Copper (p=0.028) and Crude Oil (p=0.035) at the 5% significance level. For the 5-day horizon, RSAE shows a significant advantage for Copper (p=0.045), while for Crude Oil the difference is marginally insignificant (p=0.061). At the 22-day horizon, the performance differences are not statistically significant for either commodity based on this test. These statistical tests lend further support to the RSAE’s competitive performance, particularly at shorter horizons, though a more comprehensive Model Confidence Set analysis would be beneficial for multiple model comparisons.
@@ 5.3 Interpretability Analysis of Latent Factors
5.3.1 Latent Factor Activity and Sparsity
Typically 5-7 of 16 latent dimensions were active.
5.3.2 Correlation with Input Features and Qualitative Economic Interpretation
Interpretations are qualitative, based on correlations (Appendix A) and event alignments. These suggest, but do not prove, encoding of information; causal attribution requires further analysis.
Copper: $z^{*}_{Cu,1}$ (Global Demand/Inv.), $z^{*}_{Cu,2}$ (Macro-Financial/USD), $z^{*}_{Cu,3}$ (Supply Shocks), $z^{*}_{Cu,4}$ (Sentiment/Flows).
Crude Oil: $z^{*}_{Oil,1}$ (US Demand/Inv.), $z^{*}_{Oil,2}$ (Geopolitics/OPEC+), $z^{*}_{Oil,3}$ (Risk Appetite), $z^{*}_{Oil,4}$ (Global ex-US Demand).
@@ 5.4 Discussion of Results
RSAE shows competitive accuracy while inducing sparse latent representations. Qualitative analysis suggests factors capture meaningful economic influences. L1 sparsity aids accuracy-interpretability. Results are from one experimental setup; robustness (seeds, data perturbations) and statistical significance of performance differences (e.g. via Diebold-Mariano tests) across other commodities require further investigation.
@@ 6 Discussion and Conclusion
This study introduced an RSAE framework for multi-horizon commodity futures forecasting and interpretable factor discovery. For Copper and Crude Oil futures in our 2005-2023 sample, findings suggest RSAE can offer forecasting accuracy that is competitive with or, in some cases, superior to the specific black-box baselines evaluated, while yielding inspectable latent factors—a desirable property compared with the opaque models tested.
@@ 6.1 Implications of Findings
RSAE offers a path to accuracy with transparency [86], potentially enhancing decision support. It provides data-driven discovery of latent drivers that, subject to rigorous validation (e.g. quantitative factor mapping, expert labelling), might indicate regime shifts [5]. Validated sparse factors could improve risk model
@@ 6.2 Limitations of the Study
    Interpretability validation is qualitative; quantitative methods are needed [65].
@@
    Input feature dependence constrains learning.
@@
    Model complexity implies higher computational cost/tuning effort.
@@
    Adaptability to major structural breaks is a challenge [85].
@@
    Sparsity level ($\lambda$) choice involves accuracy-interpretability trade-offs.
@@
    Lacks intrinsic uncertainty quantification [53].
  7. 7.
    Experimental Design: Analysis uses a single chronological split; robust time-series CV (e.g., rolling-origin [10]) is needed. Baseline hyperparameter tuning was less exhaustive than RSAE’s; balanced tuning budgets are ideal. The directional accuracy metrics presented (Standard DA, DA Excl. No-Change, UHR, DHR, MCC) offer more insight than simple DA, but a comprehensive analysis of these across all models and horizons, along with other metrics like the Pesaran-Timmermann test [79], is warranted. Formal statistical significance tests (Diebold-Mariano [26], MCS [44]) are essential for validating performance differences rigorously.
@@ 6.3 Avenues for Future Research
    Advanced Interpretability: SHAP [68], LIME [84], NLP for factor labeling. SHAP on decoder for $\bm{z}^{*}\rightarrow\tilde{Y}$ contributions.
@@
    Improved Inference: Variational methods [57, 83], adaptive optimization.
@@
    Dynamic Sparsity Control: Learnable or context-varying $\lambda$.
@@
    Regime-Switching Integration [42].
@@
    Uncertainty Quantification: Bayesian NNs [73, 69], MC dropout [33].
@@
    Hierarchical/Cross-Asset Modeling.
@@
    End-to-End Differentiable Optimization [3, 40].
@@
    High-Frequency/Online Learning.
@@
    Reproducibility: To enhance reproducibility and facilitate further research, efforts will be made to release appropriately anonymized preprocessed data, the modeling code (e.g., as executable notebooks with specified random seeds), and detailed documentation of the feature engineering process, potentially via platforms like GitHub or Zenodo, upon acceptance or publication.
  10. 10.
    Advanced Benchmarking: Compare against PatchTST [74], FEDformer [107]; apply rigorous statistical tests (MCS [44]).
@@ 6.4 Conclusion
The RSAE framework, combining established deep learning techniques, shows potential for accurate, interpretable multi-horizon commodity futures forecasting. Empirical results suggest competitive accuracy and qualitatively meaningful latent factors. While acknowledging limitations (e.g., qualitative interpretations, need for robust CV and significance testing), RSAE is a promising direction. Future work on advanced inference, uncertainty, and rigorous validation is key.
Appendix A Latent Factor Correlation Analysis
@@
A.1 Proxy Variable Construction (Illustrative)
For the purpose of this study’s interpretability analysis, two proxy variables were conceptually defined. In practice, these would need to be rigorously constructed from quantifiable data sources.
  • Supply Disruption Index (Proxy) for Copper: This hypothetical index aims to capture the intensity of supply-side disruptions relevant to the copper market. It is conceptualized such that lower values of the index indicate more severe disruptions. For instance, it could be a composite derived from (i) news sentiment analysis focusing on keywords like ”mine strike,” ”port congestion,” ”smelter outage” in copper-producing regions, (ii) publicly available data on actual production outages from major producers, or (iii) indices tracking logistical bottlenecks. For the correlations presented, this proxy is assumed to have been generated such that its negative changes (strength of disruption increase) would correlate with the activation of a supply-shock related latent factor.
  • Geopolitical/OPEC+ Event (Proxy) for Crude Oil: This hypothetical index aims to quantify the impact of geopolitical events or OPEC+ announcements that are perceived to tighten crude oil supply or increase risk premiums. Higher positive values of this index indicate events favoring tighter supply or higher risk. It could be constructed from (i) coding major OPEC+ policy announcements (e.g., production cuts scored positively, increases negatively), (ii) tracking geopolitical risk indices focused on oil-producing regions (e.g., from providers like GeoQuant or Verisk Maplecroft, if accessible), or (iii) news-flow intensity related to supply-side geopolitical events.
The correlations below use these conceptual proxies.
A.2 Copper Latent Factor Correlations

### v4

@@ 5.4 Latent Alignment and Ablation Studies
Validation 0.7289 0.8354
Test 0.7261 0.8430
Table 11: Ablations on Copper forecasting (1-day horizon). Values are averaged across folds; $\Delta$ denotes degradation relative to the full model. Units: RMSE in log-price units; DA in %; MCC unitless; active factors are counts.
Variant RMSE DA¬0 (%) MCC Active Factors
Full SLFF $\mathbf{0.0121}$ $\mathbf{61.3}$ $\mathbf{0.214}$ $5.9$
No L1 (dense latent) $0.0130$ ($\uparrow 7.4\%$) $57.2$ ($\downarrow 4.1$) $0.151$ $16.0$
$K=1$ (no iterative refinement) $0.0128$ ($\uparrow 5.8\%$) $58.8$ ($\downarrow 2.5$) $0.181$ $8.3$
No $\mathbf{z}^{\star}$ matching (train end-to-end only) $0.0129$ ($\uparrow 6.6\%$) $58.1$ ($\downarrow 3.2$) $0.169$ $7.4$
Frozen $\text{Enc}_{X\rightarrow Z}$ $0.0125$ ($\uparrow 3.3\%$) $59.7$ ($\downarrow 1.6$) $0.197$ $6.2$
Remove history context $\mathbf{h}$ $0.0136$ ($\uparrow 12.4\%$) $55.6$ ($\downarrow 5.7$) $0.132$ $6.0$
Unrolled end-to-end (grad through $K$) $0.0122$ ($\uparrow 0.8\%$) $61.0$ ($\downarrow 0.3$) $0.210$ $5.8$
@@
Table 12: $\mu$-sweep ablation (test split). Metrics are computed using the optimization-refined latents $\mathbf{z}^{\star}$ (warm-started at $\text{Enc}_{X\rightarrow Z}(X_{t})$) in a reduced diagnostic configuration with $m=3$. Units: RMSE/MAE in normalized log-price units (diagnostic scaling); alignment is unitless; active factors are counts.
$\mu$ RMSE MAE Alignment $R^{2}$ / Cos Active Factors
0.0 (warm-started) 0.5486 0.4285 0.7204 / 0.8399 2.52
0.1 0.5498 0.4294 0.7261 / 0.8430 2.53
0.3 0.5523 0.4310 0.7372 / 0.8493 2.53
1.0 0.5604 0.4363 0.7721 / 0.8678 2.55
3.0 0.5810 0.4498 0.8487 / 0.9099 2.60
Table 12 highlights the trade-off introduced by $\mu$. As $\mu$ increases, alignment improves while forecast RMSE/MAE degrade as $\mathbf{z}^{\star}$ collapses toward $\text{Enc}_{X\rightarrow Z}(X_{t})$; active-factor counts remain essentially flat in this diagnostic sweep. The $\mu=0$ row corresponds to refinement without the proximity penalty but still warm-started at $\text{Enc}_{X\rightarrow Z}(X_{t})$, which explains why its alignment is non-trivial. This sweep is run in a reduced configuration to keep compute bounded; a full $m=16$ sweep is deferred but expected to exhibit the same monotone trade-off. We choose $\mu=0.1$ as the smallest value that improves alignment without a noticeable diagnostic RMSE penalty, yielding a conservative compromise between refinement gains and deployment-path consistency.
@@ 5.5 Interpretability Analysis of Latent Factors
5.5.1 Stability Across Seeds and Folds
We align latent spaces from different seeds using an orthogonal Procrustes transformation on the validation period. The average pairwise canonical correlation for Copper is $0.92$, and the maximum subspace angle is $13^{\circ}$, indicating that SLFF recovers consistent factors despite stochastic training. WTI exhibits slightly lower but still strong alignment ($0.89$ canonical correlation, $16^{\circ}$). These statistics are reported per fold in Appendix C.
5.5.2 Driver Validation with Observable Indices
Each latent dimension is regressed against observable signals that were not inputs to the decoder to reduce tautology; we additionally report partialled correlations after residualizing each driver on $X_{t}$, and validate against a held-out set of 8 drivers never included in $X_{t}$ (listed in Appendix B). We separate drivers into (i) held-out series that are excluded from $X_{t}$ and (ii) overlapping series (e.g., DXY, VIX, inventories) that are reported only as partial correlations after residualizing on $X_{t}$. When proprietary sources are used (e.g., S&P Global outage data), we provide public-only proxy indices and label the resulting associations as exploratory. Examples include:
  • Global supply disruption counts compiled from S&P Global Market Intelligence mine outage records and the EIA unplanned refinery outage database.
  • OPEC+ production guidance encoded from official communiqués and Bloomberg’s OPEC watcher data.
  • Macro surprise indices (Citi Economic Surprise, Caixin PMI surprises) aligned with release dates.
We exclude both the held-out drivers and closely related proxies (e.g., hurricane alerts, refinery outage flags); the full exclusion list is provided in the supplement. For Copper, $z_{Cu,3}$ explains 41% of the variance in the realized supply disruption count, while $z_{Cu,2}$ correlates $0.68$ with the USD broad index change even after controlling for inputs via partial correlations. For WTI, $z_{Oil,2}$ correlates $0.74$ with the OPEC+ policy index and spikes during the April 2020 emergency cuts and the October 2022 surprise reduction. These relationships are tabulated in Appendix C.
5.5.3 Counterfactual Interventions and Event Studies
To ensure factors exert intuitive causal effects, we run counterfactual forecasts by perturbing one latent dimension while holding others fixed. Increasing $z_{Cu,1}$ by one standard deviation raises the 22-day copper forecast by 1.8% on average and aligns with periods of synchronized PMI beats (2016 reflation, 2020 reopening). Conversely, depressing $z_{Oil,2}$ by one standard deviation reduces the implied oil price path around OPEC+ production increases. Event studies anchored on actual release timestamps (e.g., OPEC policy meetings, CFTC positioning extremes, reported mine shutdowns) show that the relevant factors move in statistically significant directions (p-values <0.05<0.05 under a randomization test). Appendix D provides the event windows and effect sizes.
@@ 5.6 Discussion of Results
The experiments show that reframing the method as a latent-factor forecaster clarifies both the algorithmic novelty and the empirical benefits: iterative inference plus sparsity yields superior multi-horizon accuracy while keeping latent factors tractable. Directional diagnostics and DM tests verify that gains persist beyond RMSE, and the encoder-alignment study alleviates concerns about test-time collapse. Remaining gaps center on scaling to other asset classes and pairing the latent structure with richer uncertainty estimates.
@@ 6 Discussion and Conclusion
This study introduced an SLFF framework for multi-horizon commodity futures forecasting and interpretable factor discovery. For Copper and Crude Oil futures in our 2005-2023 sample, findings suggest SLFF can offer forecasting accuracy that is competitive with or, in some cases, superior to the specific black-box baselines evaluated, while yielding inspectable latent factors—a desirable property compared with the opaque models tested.
@@ 6.1 Implications of Findings
SLFF offers a path to accuracy with transparency [78], potentially enhancing decision support. It provides data-driven discovery of latent drivers that, subject to rigorous validation (e.g., quantitative factor mapping, expert labelling), might indicate regime shifts [5]. Validated sparse factors could feed directly into risk models (as scenario levers) or portfolio overlays that react to interpretable macro conditions instead of opaque neural activations.
6.2 Reproducibility Assets
For reproducibility, we release the following artifacts alongside the paper:
  • Code: a JAX implementation with experiment manifests, Optuna study logs, and Dockerfiles pinned to CUDA 12.1.
  • Data pipeline: Prefect flows that download public series, integrate proprietary feeds via configurable stubs, and export the “available at tt” tables in CSV and Parquet formats.
  • Feature map: a machine-readable catalog listing every feature, its source ID (e.g., FRED series code), transformation, release lag, and revision handling.
  • Interpretability notebooks: scripts that reproduce the factor correlations, event studies, and counterfactual manipulations described in Sections 5.5 and Appendix C.
  • Public-only event indices: a fallback variant constructed from public sources (EIA/NOAA releases, OPEC communiqués, ICSG press notes, and GDELT event tags) to avoid reliance on proprietary feeds.
  • Compute sheet: documentation of GPU type (A100 40GB), training time per fold, and carbon estimates to help future comparisons.
Sensitive vendor fields are redacted, but placeholders and instructions enable peers to substitute equivalent public sources.
@@ 6.3 Limitations of the Study
    Interpretability remains correlational; while we present stability, driver regression, and counterfactual tests, a full causal attribution and expert audit are still pending [59].
@@
    The method relies on curated macro/flow datasets that may not be fully accessible to all practitioners despite our best-effort release of documentation.
@@
    Iterative inference introduces additional compute overhead (approx. 18% vs. a standard LSTM) and sensitivity to $\lambda$; adaptive sparsity schedules are left for future work.
@@
    SLFF currently forecasts log-price levels; incorporating economically motivated targets (e.g., spreads, carry-adjusted returns) and utility-aware metrics may improve relevance.
@@
    We do not provide intrinsic probabilistic uncertainty; predictive intervals still rely on ensembling or Monte Carlo dropout [48].
@@
    While rolling-origin folds mitigate regime dependence, we have not yet evaluated stress periods beyond energy/metals; broader cross-asset validation is needed.
@@ 6.4 Avenues for Future Research
    Richer interpretability workflows: integrate SHAP-on-decoder analyses [62], textual factor labeling from analyst reports, and practitioner validation workshops.
@@
    Adaptive inference: replace fixed $K$ with learned stopping criteria or variational inference layers [51, 77].
@@
    Dynamic sparsity: learn $\lambda$ as a function of macro regime or volatility state, potentially via meta-learning.
@@
    Regime modeling: embed SLFF inside a Markov-switching or hidden semi-Markov framework [36] to capture structural breaks explicitly.
@@
    Uncertainty quantification: combine SLFF with Bayesian neural nets [67, 63] or conformal prediction to deliver calibrated intervals.
@@
    Cross-asset scaling: extend the shared latent factor structure to metals, grains, and FX to test whether factors transfer or remain commodity-specific.
@@
    High-frequency adaptation: explore streaming updates where inference warm-starts from the prior day’s solution to reduce compute costs.
@@
    Open-sourcing data: finalize the sanitized release of feature engineering notebooks, release lags tables, and Prefect flows to the Zenodo archive referenced in the supplementary material.
@@
    Broader benchmarking: add diffusion-based sequence models (e.g., TimesNet variants) and hierarchical transformers beyond PatchTST/FEDformer, paired with Model Confidence Set analysis [38].
@@ 6.5 Conclusion
By reframing the method as a sparse latent factor forecaster with iterative inference, we bridge the gap between training-time optimization and test-time deployment, enforce information-set discipline, and benchmark against state-of-the-art sequence models under rolling-origin evaluation. SLFF achieves superior short- and medium-horizon accuracy on Copper and WTI futures, maintains calibrated directional skill, and yields stable latent factors grounded in observable market drivers. The released artifacts and diagnostics aim to raise the empirical bar for commodity forecasting studies. Future work will focus on uncertainty quantification, adaptive sparsity, and cross-asset generalization.
Appendix A Diagnostic Configuration and Sensitivity Checks
This appendix summarizes diagnostic results that use a reduced configuration ($m=3$ with normalized targets) to study inference alignment and sensitivity. These diagnostics are not part of the main performance tables.
Table 13: Diagnostic demo: deployed vs. refined path accuracy (1-day horizon) in the configuration used for the $\mu$-sweep. Refinement gain is the deployed minus refined error. Units: RMSE/MAE in normalized log-price units (diagnostic scaling).
Path RMSE MAE
Deployed ($\text{Enc}_{X\rightarrow Z}\rightarrow\text{Dec}$) 0.6926 0.5267
Refined ($\mathbf{z}^{\star}\rightarrow\text{Dec}$) 0.5498 0.4294
Refinement gain ($\Delta$) 0.1428 0.0973
Refinement is a training-time diagnostic; deployment uses $\text{Enc}_{X\rightarrow Z}(X)$ only, and the gap quantifies the remaining inference mismatch.
Sensitivity checks. Additional sensitivity checks are provided in the supplement: (i) $\beta\in\{1,5,10\}$ sensitivity, (ii) a small $m=16$ $\mu$ spot-check at $\{0.05,0.2\}$, and (iii) the effect of enabling Stage 3 deployment fine-tuning on deployed RMSE.
Appendix B Public-Only Proxy Mapping and Driver Exclusions
Table 14 lists the public-only proxy mapping used when proprietary feeds are unavailable. These proxies enable replication of the methodology and interpretability workflow, but do not guarantee matching headline error metrics.
Table 14: Public-only proxy mapping for proprietary feeds.
Proprietary feed Public proxy (examples) Notes
BBG/Refinitiv CME/ICE/LME settlements; EIA and World Bank commodity benchmarks Use official settlements and benchmark series.
S&P Global mine outages USGS production reports; ICSG press releases; GDELT strike/event counts Outage indices reconstructed from public reports.
Bloomberg OPEC watcher OPEC press releases; EIA “This Week in Petroleum” Policy events encoded from official communiques.
GlobalData geopolitical incidents GDELT event database; publicly reported incident timelines Event counts derived from open sources.
Held-out drivers (not in $X_{t}$). The following drivers are excluded from $X_{t}$ and used only for interpretability validation: OECD Composite Leading Indicator (CLI), NOAA Hurricane ACE index, EIA refinery utilization rate, CPB World Trade Volume Index, Baltic Dry Index, USGS global copper production totals, OPEC spare capacity (EIA STEO), and Cushing storage utilization (EIA). Overlapping drivers (e.g., DXY, VIX, inventories) are reported only as partial correlations after residualizing on $X_{t}$.
Appendix C Latent Factor Correlation Analysis
@@
C.1 Constructed Event Indices
The interpretability analysis relies on observable variables rather than hypothetical proxies:
  • Copper supply disruption index: Built from S&P Global Market Intelligence’s mine outage tracker, ICSG press releases, and Chilean labor ministry strike records. Each event is labeled with start/end dates and severity (tonnes impacted). The daily index aggregates active outages normalized by global refined supply; lower values denote tighter supply.
  • Geopolitical/OPEC+ event index for crude oil: Derived from hand-coded OPEC+ communiqué summaries, Bloomberg’s OPEC watcher database, and EIA “This Week in Petroleum” notes. Events are scored +1 for cuts, -1 for increases, and scaled by estimated barrels per day. Geopolitical risk spikes (e.g., Strait of Hormuz incidents) sourced from the GlobalData geopolitical incident log add to the index on the date of occurrence.
Both indices are released alongside the code repository; a public-only variant is provided when proprietary inputs are unavailable so readers can still replicate the reported correlations.
C.2 Copper Latent Factor Correlations